# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
$A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
$(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.6 (9: $v\neq 1$, gen, $U(1)_{1}=\infty$)
@@
If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive light weight with $v\in V(h)$ and if $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h,v})_{1/\xi}(1)_{1}=+\infty\,.$
@@
Proof. We denote the minimum length of a walk joining $1$ and $v$ in $G(h)$ by $m$. The inner vertices of it differ from $v$. By the subseries bound principle we get $\xi^{-m}c^{h,v}_{m}>0$. By item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$ and we conclude that $\xi^{-n-m}a_{n+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\xi^{-n}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-j-m}a_{j+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
for $n\to\infty$, and $(A_{h,v})_{1/\xi}(1)_{1}=+\infty$. $\Box$
We proceed to $\xi$-convex weights.
Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
@@
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and $(C_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,.$}
@@
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$ and $C_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,.$}
Proof. 1. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 4.1 we have $(C_{h,v})_{1/\xi}(x)^{2}=1-\frac{(B_{h,v})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22, $\big((C_{h,v})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}$. The stated formula follows.
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
Again, if we replace the assumption of non-vanishing of $B_{h,v}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
Theorem 4.8 (10: $v\neq 1$, con, $U(1)_{1}$)
@@
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{h,v})_{1/\xi}(1)_{1}$ and $(E_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,$,}
where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{h,v})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
@@
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h,v}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,$,}
@@
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. 1. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, we use the previous proof. If $(B_{h})_{1/\xi}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.20 and 2.19 that also $(B_{h,v})_{1/\xi}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $(C_{h,v})_{\xi}(1)_{1}=0$. Thus, as in the previous theorem,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=0\,.$
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.10.
Theorem 4.9 (11: $v\neq 1$, con, $U(1)_{2}$)
@@
1. Suppose that $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
@@
and that $(B_{h,v})_{1/\xi}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
@@
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}}\big)\,.$
@@
2. Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
@@
and that $B_{h,v}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $(B_{h})(1)_{2}\neq 0$ and
$A_{h,v}(1)_{2}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
@@
1. Let $h$ be $\xi$-converx with $\xi\in\mathbb{C}_{1}$, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}={\textstyle(C_{h,\,v})_{1/\xi}(1)_{1}=(C_{h,\,v})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
@@
2. Let $h$ be 0-convex, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. The beginning of the previous case gives for $\xi=1$ that $\big(C_{h,v}(1)_{2}\big)^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)$, the stated formula follows. $\Box$
Theorem 4.10 (12: $v\neq 1$, con, $U(1)_{1}=\infty$)
@@
1. If $\xi\in\mathbb{C}_{1}$, $v$ is in $\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=1\,.$

### v8

@@ 4.2 Pólya’s theorems for $v\neq 1$ and general weights
$A_{v,\,h}(1)_{2}\in D_{h}(1)_{2}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
Proof. If $B_{v,h}(x)=B_{h}(x)$, then by item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$. Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
@@
Let $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
$\big(C_{v,\,h}(1)_{2}\big)^{2}=1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{v,h}(1)_{2}=C_{v,h}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.8 ($U(1)_{1}=\infty$)
@@
Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{v,\,h})_{1/\xi}(1)_{1}=+\infty\,.$
@@
Proof. Let $m$ be the minimum length of a walk $w$ joining $1$ and $v$ in $G(h)$. Then $\xi^{-m}h(w)>0$ and $w\in C_{v,m}$. Using the subseries bound we get $\xi^{-m}h(C_{v,m})>0$. By item 1 of Proposition 4.2 we have $h(A_{v,n})=\sum_{j=0}^{n}h(C_{v,j})h(D_{n-j})$ and we conclude that
$\xi^{-n-m}h(A_{v,\,n+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\xi^{-n}h(D_{n})$
for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-j-m}h(A_{v,\,j+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\sum_{j=0}^{n}\xi^{-j}h(D_{j})\to+\infty$}
for $n\to\infty$. So $(A_{v,h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
4.3 Pólya’s theorems for $v\neq 1$ and $\xi$-convex weights
We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.3.
Theorem 4.9 ($\exists\,U(1)_{1}$ and $h\geq 0$)
@@ 4.3 Pólya’s theorems for $v\neq 1$ and $\xi$-convex weights
Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{h})_{1/\xi}(1)_{1}$ exists, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=\sqrt{1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}}\in(0,\,1)\,.$
Proof. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. Since $B_{n},B_{v,n},C_{v,n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that $(B_{h})_{1/\xi}(x)$, $(B_{v,h})_{1/\xi}(x)$ and $(C_{v,h})_{1/\xi}(x)$ have all coefficients in $[0,1]$ and are in $H_{1}$. Since $B_{v,n}\subset B_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that
$0\leq(B_{v,\,h})_{1/\xi}(1)_{1}\leq(B_{h})_{1/\xi}(1)_{1}<+\infty\,.$
By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. Using the arithmetic of functional limits, Proposition 2.20 (more precisely, the version for nonnegative coefficients), the bound $(B_{h})_{1/\xi}(1)_{1}\geq 1>0$ and Theorem 2.18, we compute that
$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=((C_{v,\,h})_{1/\xi})^{2}(1)_{1}$ $=$ $\lim_{x\nearrow 1}((C_{v,\,h})_{1/\xi})^{2}(x)$
$=$ ${\textstyle 1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}}$
$=$ ${\textstyle 1-\frac{(B_{v,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\ \ (\in[0,\,1])\,.}$
We used the third Abelian theorem on the first and third lines. Thus, due to nonnegativity, $(C_{v,\,h})_{1/\xi}(1)_{1}$ equals to the displayed root. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives that
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}$
and the stated formula follows. The membership to $(0,1)$ follows from the inequality
$1\leq(B_{v,\,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}\,.$
The strict inequality is witnessed by any walk in $G(h)$ from $1$ to $v$, which yields (by going back via the reversal) a walk in $G(h)$ from $1$ to $1$ visiting $v$. $\Box$
Theorem 4.9 covers only the finite case $(B_{h})_{1/\xi}(1)_{1}<+\infty$; we move the case $(B_{h})_{1/\xi}(1)_{1}=+\infty$ to part 1 of Theorem 4.13.
Theorem 4.10 ($\exists\,U(1)_{1}$)
@@
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{v,h})_{1/\xi}(1)_{1}$ exist and $(B_{h})_{1/\xi}(1)_{1}\neq 0$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,.$
@@
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$ and $C_{v,h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
Proof. 1. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22,
$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\,.$}
The stated formula follows.
2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. We use the displayed formula for $(C_{v,h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
We again trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
Theorem 4.11 ($\exists\,U(1)_{1}$)
@@
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{v,h})_{1/\xi}(1)_{1}$ and $(E_{v,h})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,$,
@@
where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{v,h})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
@@
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
Proof. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, resp. $B_{h}(1)_{1}\neq 0$, we use the previous proofs. We therefore assume that $(B_{h})_{1/\xi}(1)_{1}=0$, resp. $B_{h}(1)_{1}=0$.
1. From the proof of part 1 in Theorem 4.10 we know that
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
Using item 2 of Proposition 4.2 and Propositions 2.20 and 2.19, it follows from $(B_{h})_{1/\xi}(1)_{1}=0$ that $(B_{v,h})_{1/\xi}(1)_{1}=0$. Item 3 of Proposition 4.2 and Proposition 2.20 give $(C_{v,h})_{1/\xi}(1)_{1}=0$. So the displayed limit is 0.
2. Since $\xi=0$, item 1 of Proposition 4.2 and Proposition 2.31 give that $A_{v,h}(x)=C_{v,h}(x)$. As we know, $B_{h}(1)_{1}=0$ implies that $B_{v,h}(1)_{1}=0$ and that $C_{v,h}(1)_{1}=0$. So $A_{v,h}(1)_{1}=0$. $\Box$
The last two theorems generalize the case $d\geq 3$ of the formula (1) for $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.13.
Theorem 4.12 ($\exists\,U(1)_{2}$)
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
@@
and if $(B_{v,h})_{1/\xi}(1)_{2}$ exists, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
@@
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\Big)\,.$
@@
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
@@
and if $B_{v,h}(1)_{2}$ exists, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
then $B_{h}(1)_{2}\neq 0$ and
$A_{v,\,h}(1)_{2}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
Proof. Let $B_{v,h}(x)=B_{h}(x)$. By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$. Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
@@
1. Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\big((C_{v,h})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}\xi^{-n}h(A_{v,n})={\textstyle(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
@@
2. Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By part 1 for $\xi=1$ we have $\big(C_{v,h}(1)_{2}\big)^{2}=1-\frac{B_{v,h}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)$. The stated formula follows. $\Box$
Theorem 4.13 ($U(1)_{1}=\infty$)
@@
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=1\,.$